MICROWAVE SPECTRUM AND STRUCTURE DETERMINATION OF THE CCP AND ^{13}C^{13}CP RADI-CALS (X^{2}Π_{r})

\underline{M. SUN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721; D. T. HALFEN, Steward Observatory, University of Arizona, Tucson, AZ, 85721; D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY, 40506; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The microwave spectrum of the CCP and ^{13}C^{13}CP radicals (X^{2}Π_{r}) has been measured using Fourier Transform techniques. These species were created by the reaction of a mixture of PCl_{3} vapor and acetylene diulted in argon carrier gas and a DC glow discharge. Two rotational transitions each were measured in the frequency range of 18 to 32 GHz. The lambda-doublets of the Ω = 1/2 component of CCP and ^{13}C^{13}CP were each split by phosphorus hyperfine interactions. For ^{13}C^{13}CP, additional hyperfine splittings were observed due to both ^{13}C atoms, creating several doublets of doublets. These data were analyzed with a Hamiltonian incorporating three nuclear spins. The data were combined with millimeter/submillimeter measurements for CCP and ^{13}C^{13}CP, and fit with a case (c) Hamiltonian, and effective rotational, lambda-doubling, and phosphorus and carbon-13 hyperfine constants were determined.